OBSERVATIONS OF H_{3}^{+} IN THE DIFFUSE INTERSTELLAR MEDIUM

\underline{B. J. McCALL}, Department of Astronomy & Department of Chemistry, University of California at Berkeley, 601 Campbell Hall, Berkeley, CA 94720; T. R. GEBALLE, Gemini Observatory, Hilo, HI 96720; K. H. HINKLE, National Optical Astronomy Observatories, Tucson, AZ 85726; M. GOTO, N. KOBAYASHI, H. TERADA, T. USUDA, National Astronomical Observatory of Japan, Hilo, HI 96720; T. OKA, Department of Chemistry and Department of Astronomy & Astrophysics, University of Chicago, Chicago, IL 60637.

The unexpectedly high column density of H_{3}^{+} observed^{a}^{,}^{b} in the diffuse interstellar medium towards Cygnus OB2 12 presented quite an enigma for interstellar chemistry. Since H_{3}^{+} has now been detected in many sightlines^{c}, it is clear that there is a general problem with the chemical models.

The standard model of H_{3}^{+} chemistry in diffuse clouds contains only three parameters: ζ (the H_{2} ionization rate), k_{e} (the H_{3}^{+} dissociative recombination rate constant), and [e]/[H_{2}] (the electron fraction) — evidently, at least one of the assumed values for these parameters is in error by one to two orders of magnitude. There are only three options: (1) the electron fraction is lower than expected from complete conversion of C → C^{+}, (2) the value of k_{e} for J = 1 H_{3}^{+} (and cold electrons) is lower than that reported in the laboratory for rotationally hot H_{3}^{+}, and/or (3) the value of ζ is higher in diffuse clouds than in dense clouds due to the presence of low-energy cosmic rays.

We report new observations of H_{3}^{+} in the diffuse interstellar medium using CGS4 at UKIRT and IRCS at Subaru, which seem to rule out option (1). In particular, we have tentatively detected H_{3}^{+} in the classical diffuse cloud towards ζ Per, toward which both C^{+} and H_{2} have been spectroscopically measured.

It now seems that it is up to the dissociative recombination community to unambiguously determine the appropriate value of k_{e} through both experiment and theoretical calculation. With a definitive value of k_{e} in hand, observations of H_{3}^{+} will then serve as a direct probe of the cosmic-ray ionization rate ζ in diffuse clouds.

aB. J. McCall, T. R. Geballe, K. H. Hinkle, and T. Oka, Science 279, 1910 (1998)

bT. R. Geballe, B. J. McCall, K. H. Hinkle, and T. Oka, Astrophysical Journal 522, 338 (1999)

cB. J. McCall et al. Astrophysical Journal 567, 391 (2002)